OpenStudy (anonymous):

Compute the higher order derivative : f(t) =tcost, find the f ' ' ' ( 0 )

OpenStudy (anonymous):

Anyone please solve and explain .

OpenStudy (anonymous):

we need to find laplace derivatives?

OpenStudy (anonymous):

Then ?

OpenStudy (anonymous):

What do you mean Laplace derivative . ? heres another example \[g(x) = \sqrt{5-2x} , (Find ) g ' ' ' (2)\]

OpenStudy (anonymous):

\[L(\cos(t))=S/S^2+1\]? in this form?

OpenStudy (sidsiddhartha):

theren is no need of laplace here @aryandecoolest @EinsteinMorse just use product rule take u=t v=cost

OpenStudy (anonymous):

ohk!! @sidsiddhartha \[dy/dx= u(dv/dx)+v(du/dx)\] ------> use this rule

OpenStudy (sidsiddhartha):

yeah that'll be enough :)

OpenStudy (anonymous):

But its say that Compute the Higher Order Derivative of That ?

OpenStudy (sidsiddhartha):

yeah just keep on differentiating 3 times as it says f'''(x) after that put x=0 and u'll get f'''(0)

OpenStudy (anonymous):

yeah at end after differentiating 3 times, put x=0

OpenStudy (anonymous):

Then What is the meaning of f ( 0 ) ?

OpenStudy (anonymous):

Then in the answer is -3 ?

OpenStudy (anonymous):

substitute x=0

OpenStudy (anonymous):

I got the derivative 3 times then I substitute the 0 to t , Then I got -3

OpenStudy (anonymous):

good. is that correct?. If your procedure is right it should be.

OpenStudy (anonymous):

What about the Laplace Derivative . ? im curious about it :)

OpenStudy (anonymous):

you know what laplace derivatives are? @EinsteinMorse

OpenStudy (anonymous):

I dont know The Laplace derivative and also The partial derivative ?

OpenStudy (anonymous):

In this differentiation and integration being performed in the "t"-space are transformed these expressions into the "s"-space expressions...

OpenStudy (anonymous):

so if you wanted to calculate tcost from that then \[L(tcost)=(1/2(1/i-s^2) +1/(i+s^2))\] \[s^2-1/(s^2+1)^2\]

OpenStudy (anonymous):

Thats the Laplace

OpenStudy (hysenberg):

@aryandecoolest there's a bit mistake in the laplace it will be \[\frac{ 1-s^2 }{ (1+s^2)^2 }\]

OpenStudy (anonymous):

is it @hysenberg ? ) anyways thanks for the verification

OpenStudy (anonymous):

What is the Partial Derivative ?

OpenStudy (anonymous):

\[d^2f/dx^2= f xx\] \[d^2f/dxdy=fxy\] \[d^3x/d^2x dy= fx xy\]

OpenStudy (anonymous):

Hows that formula Works :)